# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v2

@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 63.37 52.19 69.16 60.40 64.36 60.78
Score-level Average 52.00 50.50 58.42 51.56 57.94 61.39 60.40 56.03
Concat 50.46 59.41 62.38 52.19 60.75 56.44 57.43 57.01
SE-Concat 55.38 48.51 58.42 52.81 62.62 57.43 59.41 56.37
Cross-Atten 52.31 53.47 58.42 51.25 64.49 58.42 57.43 56.54
MLP-Mixer 53.23 60.40 62.38 52.19 61.68 58.42 58.42 58.10
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 63.37 53.75 58.88 64.36 63.37 60.56
Score-level Average 48.62 59.41 51.49 51.25 60.75 62.38 63.37 56.75
Concat 57.23 60.40 56.44 51.25 56.07 57.43 56.44 56.47
SE-Concat 55.08 60.40 59.41 54.06 60.75 61.39 57.43 58.36
Cross-Atten 50.15 56.44 61.49 57.50 59.81 64.36 58.42 58.31
MLP-Mixer 52.62 56.44 58.42 52.50 60.75 61.39 58.42 57.22
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 57.43 54.06 66.36 64.36 60.40 58.68
Score-level Average 53.85 51.49 65.35 53.75 58.88 57.43 58.42 57.02
Concat 52.92 52.48 55.45 53.75 62.62 62.38 56.44 56.58
SE-Concat 51.38 60.40 58.42 51.56 57.01 62.38 59.41 57.22
Cross-Atten 51.80 52.48 62.38 53.12 59.81 64.36 60.40 57.76
MLP-Mixer 53.85 55.45 60.40 50.94 59.81 64.36 60.40 57.89
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 59.41 55.00 58.88 59.41 60.40 58.60
@@ 4.5 Cross-domain Testing with Multimodal Fusion
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [50], Cross-Atten [44], MLP-Mixer [21], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [21] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). It results in four combinations of either two or three types of inputs. As shown in Table IV, among all the fusion methods, the proposed Attention-Mixer fusion (Atten-Mixer) achieved the best results on average.
Multi-to-Single Domain with Fusion. We benchmarked multi-to-single cross-domain generalization with different fusion methods by three types of cross-testing strategies. In total, seven sub-experiments were conducted for different domain combinations. The results are shown in Tabel V. On average, the best accuracies were 58.43%, 59.13%, and 60.78% on domain-simultaneous, domain-alternating, and domain-by-domain strategies, respectively. Among these, for both domain-simultaneous and domain-by-domain strategies, the best results were achieved by taking visual (face frames) and visual (AU+gaze+affect) features as input, while the best result for domain-alternating was achieved by using visual (face frames), visual (AU+gaze+affect) and audio (Mel spectrogram) features. Taking a close look at the average fusion results, for each type of modality input, Atten-Mixer achieved the best among the six fusion methods. showing the effectiveness of the proposed method. By using Atten-Mixer, in general, the average result showed slightly better when using the domain-by-domain strategy and taking visual (face frames) and visual (AU+gaze+affect) features as input. To sum up, the results showed that on the current publicly available datasets, visual features were better when it came to multi-to-single cross-domain generalizability. However, the performance differences were not a large gap, and there were also no significant differences by comparing two-to-one domain and three-to-one domain cross-testing performances.
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table IV from the top to the bottom.
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table IV are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers.
@@
TABLE VI: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM). The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 Avg
Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 58.43
IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 57.90
Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 59.69 (+1.26)
Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 58.14
IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 58.02
Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 58.26 (+0.12)
Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 56.78
IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 55.97
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 57.23 (+0.45)
Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.62
IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 57.77
@@
Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 57.43 61.39 55.57 59.46 62.58 60.60 58.54 (+0.92)
Performance of MM-IDGM We proposed MM-IDGM and conducted experiments for multi-to-single cross-domain generalization, with the results presented in Table VI. Both the IDGM and MM-IDGM algorithms are developed based on the domain simultaneous sampling method. Using the proposed Atten-Mixer fusion method as the baseline, we applied both the original IDGM and the proposed MM-IDGM. The results show that MM-IDGM consistently improves average accuracies across different modalities and inputs. However, cross-domain generalization in deception detection remains highly challenging, as some accuracies even decreased after applying IDGM. This may be due to the difficulty in optimizing the inner product between different domains. While MM-IDGM may not always perform optimally in certain cases, such as testing on the B2 dataset where dataset quality affects the results, it generally shows better performance in cross-domain generalization.
@@ 4.6 Network Details
Here, we present the detailed network architectures in Table VII, including encoders, the fusion module, and classifiers. This network architecture is utilized for experiments on multi-to-single cross-domain generalization with fusion.
TABLE VII: Architecture details for different modalities and fusion module.

### v3

@@ 4.4 Domain-Simultaneous with Gradient Reversal Layer (GRL)
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 57.54 58.42 53.62 54.32 63.37 52.19 58.04 53.78 69.16 60.40 64.36
Score-level Average 52.00 50.50 50.32 50.87 58.42 51.56 56.98 50.42 57.94 61.39 60.40
Concat 50.46 59.41 51.24 50.95 62.38 52.19 57.25 51.28 60.75 56.44 57.43
SE-Concat 55.38 48.51 51.35 52.54 58.42 52.81 57.83 51.87 62.62 57.43 59.41
Cross-Atten 52.31 53.47 52.37 54.68 58.42 51.25 57.39 52.33 64.49 58.42 57.43
MLP-Mixer 53.23 60.40 52.45 54.08 62.38 52.19 57.89 52.39 61.68 58.42 58.42
@@
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 53.85 66.34 54.32 54.89 63.37 53.75 57.96 53.04 58.88 64.36 63.37
Score-level Average 48.62 59.41 51.28 50.88 51.49 51.25 56.83 51.28 60.75 62.38 63.37
Concat 57.23 60.40 51.35 50.39 56.44 51.25 57.38 51.38 56.07 57.43 56.44
SE-Concat 55.08 60.40 53.86 52.83 59.41 54.06 57.92 52.98 60.75 61.39 57.43
Cross-Atten 50.15 56.44 54.89 53.58 61.49 57.50 57.85 52.96 59.81 64.36 58.42
MLP-Mixer 52.62 56.44 54.78 53.86 58.42 52.50 58.96 53.87 60.75 61.39 58.42
@@
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 51.69 56.44 55.06 54.87 57.43 54.06 58.91 53.48 66.36 64.36 60.40
Score-level Average 53.85 51.49 54.76 50.38 65.35 53.75 55.64 50.32 58.88 57.43 58.42
Concat 52.92 52.48 55.93 51.77 55.45 53.75 55.84 51.28 62.62 62.38 56.44
SE-Concat 51.38 60.40 56.87 52.36 58.42 51.56 56.92 51.98 57.01 62.38 59.41
Cross-Atten 51.80 52.48 56.82 53.58 62.38 53.12 57.83 52.38 59.81 64.36 60.40
MLP-Mixer 53.85 55.45 57.82 53.69 60.40 50.94 58.92 53.56 59.81 64.36 60.40
@@
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 55.69 61.39 58.05 55.72 59.41 55.00 60.31 54.34 58.88 59.41 60.40
Modality & Inputs Fusion Postion Fusion Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Score-level Average 50.76 53.62 52.31 51.22 49.62 50.38 52.35 51.15 54.37 52.12 52.98
Concat 51.25 54.21 52.38 51.42 53.21 52.46 53.46 52.31 54.88 54.38 54.49
SE-Concat 51.36 54.38 52.36 52.36 54.87 53.49 54.63 52.33 55.62 55.72 54.27
Cross-Atten 52.45 55.32 53.67 53.34 55.74 54.42 56.67 53.87 55.71 55.84 55.45
MLP-Mixer 52.66 55.48 53.48 53.76 56.72 56.73 56.72 54.83 56.82 55.71 56.78
Visual (Face frames) + Visual (AU+Gaze+Affect) Feature-level Atten-Mixer(Ours) 53.86 57.34 53.86 54.11 58.24 57.02 57.39 54.87 58.78 58.72 57.59
Score-level Average 50.38 53.25 52.33 52.39 54.42 51.25 52.38 51.24 54.31 53.43 53.63
Concat 51.49 53.48 52.47 53.18 56.31 53.32 54.32 52.33 54.33 56.72 54.68
SE-Concat 52.95 55.28 52.38 53.91 57.82 54.75 55.64 54.32 54.21 56.83 55.06
Cross-Atten 52.86 56.83 53.34 53.78 58.62 57.33 55.87 54.65 55.53 56.79 55.61
MLP-Mixer 53.47 56.74 53.48 54.27 58.72 57.38 57.64 54.87 55.49 57.83 56.35
Visual (Face frames) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.87 58.69 55.87 54.86 58.98 58.24 58.99 55.68 61.39 58.99 58.13
Score-level Average 52.86 56.31 51.85 52.38 53.21 51.27 57.67 51.23 54.06 54.72 54.43
Concat 52.65 55.89 51.64 53.34 56.72 52.23 57.83 52.35 54.65 55.49 54.69
SE-Concat 53.49 57.28 52.38 54.96 57.46 56.93 58.72 52.39 55.68 57.89 56.35
Cross-Atten 53.64 58.72 53.03 54.73 57.68 56.73 58.31 53.48 57.83 59.83 56.73
MLP-Mixer 54.67 59.81 53.48 55.92 58.92 57.84 59.98 54.63 58.75 59.86 56.95
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 54.78 60.45 54.72 56.03 59.21 58.82 60.12 55.69 60.33 62.31 57.89
Score-level Average 52.38 57.33 52.88 53.42 56.33 52.31 54.32 51.25 53.46 58.93 54.90
Concat 53.08 58.31 53.02 54.33 56.31 53.42 54.76 52.38 54.55 58.72 55.23
SE-Concat 54.89 58.39 53.76 55.41 57.35 56.65 55.65 54.83 56.73 58.79 56.24
Cross-Atten 55.98 59.74 53.47 55.64 57.36 57.82 57.86 55.73 56.79 58.22 56.84
MLP-Mixer 56.73 60.28 55.08 58.02 58.72 58.93 59.93 55.76 58.32 59.89 57.66
Visual (Face frames) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) Feature-level Atten-Mixer(Ours) 57.47 63.48 56.71 58.36 59.55 60.33 61.35 57.83 60.36 62.35 58.88
@@ 4.5 Cross-domain Testing with Multimodal Fusion
Single-to-Single Domain with Fusion. In this section, we conducted experiments of single-to-single cross-domain testing. Two types of fusion positions were involved including score-level and feature-level fusion. For feature-level fusion, multiple fusion methods were adopted such as simple concatenation, SE-Concat [55], Cross-Atten [47], MLP-Mixer [26], and Attention-Mixer fusion (Ours). To be specific, simple concatenation refers to the concatenation of the extracted features before the input to the classifier. SE-concat stands for the SE attention applied to the concatenated features. Cross-Atten means the crossmodal attention among input features by using the attention mechanism from the Transformer. MLP-Mixer uses the method in  [26] on the extracted features. The modalities and input include three types, visual (face frames), visual (AU+gaze+affect), and audio (Mel spectrogram). As shown in Table V, there are 25 domain transfer tasks, and the proposed Atten-Mixer consistently outperforms other methods across all modalities and fusion types, achieving the highest average accuracies. This demonstrates its strong ability to learn transferable representations. Feature-level fusion generally yields better performance than score-level fusion, highlighting the benefit of early interaction between modalities. Among the modalities, the best average performance across all the datasets is achieved when fused with Atten-Mixer.
Multi-to-Single Domain with Fusion. Tables VIVIII present multi-to-single domain generalization results across 21 transfer tasks using four different modality combinations under three sampling strategies. Across all settings, the proposed Atten-Mixer consistently outperforms baseline fusion methods, achieving the highest average accuracies. Notably, feature-level fusion of Visual (AU+Gaze+Affect) + Audio using Atten-Mixer reaches the best overall average performance: 57.97% (Simultaneous), 56.35% (Alternating), and 58.88% (domain-by-domain), showing its strong cross-domain generalization. Among different strategies, domain-by-domain consistently yields the highest averages across modalities, similar to the observation in the multi-to-single domain generalization results without fusion methods. Besides, feature-level fusion generally outperforms score-level fusion, indicating the importance of deep modality interaction for fusion. Additionally, multimodal combinations (e.g., Visual + Audio) consistently outperform unimodal inputs, validating the effectiveness of multimodal fusion methods on visual and audio cues in deception detection. Overall, the results show the effectiveness of Atten-Mixer and underscore the advantage of modality fusion in audio-visual deception detection.
@@
TABLE IX: The fusion results of multi-to-single cross-domain generalization accuracy (%) for multimodal inter-domain gradient matching (MM-IDGM). The numbers in the brackets show the improved accuracies compared with Atten-Mixer.
Modality & Inputs Method R&M to B1 R&M to B2 R&B1 to B2 R&B1 to M B1&M to R B1&M to B2 R&B1&M to B2 R&M to D R&M to E R&B1 to D R&B1 to E
Atten-Mixer(Ours) 53.23 57.43 63.37 52.19 57.01 63.37 62.38 53.50 53.40 57.39 51.86
IDGM 55.22 57.45 59.85 51.41 61.34 59.41 60.61 52.70 52.99 57.05 51.22
Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 57.85 58.42 60.40 52.50 65.22 62.05 61.42 53.88 54.02 58.51 52.87
Atten-Mixer(Ours) 55.69 64.36 56.44 53.75 58.88 59.41 58.42 53.97 53.43 57.82 52.35
IDGM 54.43 60.40 61.07 51.40 59.21 59.88 59.75 53.02 53.11 56.24 52.36
Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.54 59.41 62.38 51.56 59.07 62.38 59.50 54.26 54.29 58.33 53.46
Atten-Mixer(Ours) 55.08 55.45 58.42 54.37 52.34 61.39 60.40 57.39 54.20 58.02 52.83
IDGM 55.49 54.92 58.36 50.31 53.15 60.35 59.21 56.87 53.40 57.54 52.33
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 56.05 56.03 59.37 52.41 53.40 62.42 60.98 58.24 55.06 58.72 51.24
Atten-Mixer(Ours) 52.31 59.41 58.42 52.50 58.88 61.39 60.40 57.30 54.57 59.41 52.84
IDGM 52.45 57.02 60.42 54.25 58.42 61.79 60.03 58.02 54.21 59.45 53.14
@@
Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 52.77 58.43 61.39 55.57 59.46 62.58 60.60 58.75 55.87 60.85 53.01
Modality & Inputs Method D&E to B1 D&E to B2 D&E to M D&E to R R&B1&M to D R&B1&M to E R&D&E to B1 R&D&E to M R&M&D&E to B2 R&B1&M&D&E to B2 Average
Atten-Mixer(Ours) 53.47 59.01 53.62 54.21 58.43 57.10 57.54 53.75 60.40 60.99 56.84
IDGM 54.12 58.04 54.36 58.92 59.03 57.84 57.93 54.46 60.24 60.54 56.89
Visual (Face frames) + Visual (AU+Gaze+Affect) MM-IDGM 54.88 59.26 54.38 59.34 58.28 58.28 58.66 55.25 60.34 60.58 57.92 (+1.08)
Atten-Mixer(Ours) 54.24 59.41 53.50 55.14 58.53 58.65 58.92 54.88 62.38 59.35 57.12
IDGM 53.29 59.88 54.01 55.39 58.98 59.35 58.43 55.34 62.57 60.42 57.07
Visual (Face frames) + Audio (Mel spectrogram) MM-IDGM 53.56 59.81 55.36 56.03 59.21 59.32 59.42 55.23 61.92 62.35 57.64 (+0.52)
Atten-Mixer(Ours) 55.11 62.38 54.02 56.28 58.48 57.65 60.19 54.64 60.50 60.13 57.11
IDGM 55.87 61.24 55.02 56.21 57.03 57.95 60.25 55.32 60.61 59.85 56.73
Visual (AU+Gaze+Affect) + Audio (Mel spectrogram) MM-IDGM 55.99 62.68 55.36 57.02 59.31 58.34 60.58 55.41 60.89 61.78 57.68 (+0.57)
Atten-Mixer(Ours) 57.48 63.17 54.35 58.88 59.15 58.96 60.42 54.56 62.58 60.42 57.97
IDGM 56.98 62.11 54.68 58.25 59.64 59.31 60.15 54.73 61.99 61.23 58.01
Visual (Face frames)+ Visual (AU+Gaze+Affect)+ Audio (Mel spectrogram) MM-IDGM 58.13 64.22 55.05 59.31 60.14 59.46 61.88 56.93 62.64 62.34 59.02 (+1.05)
Performance of MM-IDGM We proposed MM-IDGM to enhance multi-to-single domain generalization in deception detection. Table IX presents the experimental results under the domain-simultaneous sampling strategy, comparing Atten-Mixer (baseline), IDGM, and the proposed MM-IDGM across various modality combinations. In all tested configurations, MM-IDGM consistently outperforms Atten-Mixer, achieving up to +1.08% improvement in average accuracy, particularly in the case of Visual (Face Frames) + Visual (AU+Gaze+Affect). This demonstrates the effectiveness of aligning gradient directions across domains at the feature level in a multimodal setting.
While IDGM improves performance in many cases, it occasionally degrades accuracy. This may be due to conflicts from aligning gradients independently per modality. In contrast, MM-IDGM aligns all modalities jointly, yielding more stable and transferable features, and showing promising improvements in multimodal cross-domain generalization. However, its effectiveness can vary depending on the complexity of the target domain, data quality, and the severity of distribution shifts.
Figure 4: Ablation study for attention-mixer layers. The number of layers 4, 5, 6, and 7 are compared. The modality and inputs for A, B, C, and D are in line with those in Table V from the top to the bottom.
Ablation Study for Attention-Mixer Fusion Module. We conducted an ablation study for the proposed attention-mixer fusion module with the changes in the number of attention-mixer layers. The experiments were conducted on single-to-single domain testing, where the average accuracies were compared. As shown in Fig. 4, the number of attention-mixer layers was set to 4, 5, 6, and 7. The modalities and inputs in Table V are compared, where “A” had the inputs of Visual (Face frame) + Visual(AU+Gaze+Affect), “B” had Visual (Face frames) + Audio(Mel spectrogram), “C” had Visual (AU+Gaze+Affect) + Audio (Mel spectrogram), and “D” had Visual (Face frame) + Visual (AU+Gaze+Affect) + Audio (Mel spectrogram). The results showed that models with 6 attention-mixer layers achieved the best average accuracies, followed by 7 attention-mixer layers. The results showed that deeper fusion networks (such as 6 layers vs. 4 or 5 layers) improve generalization by enhancing cross-modal interactions, but overly deep architectures (e.g., 7 layers) may bring diminishing returns. Therefore, it is important to carefully select the fusion depth to balance representation capacity and the risk of overfitting.
@@ 4.6 Network Details
Here, we present the detailed network architectures in Table X, including encoders, the fusion module, and classifiers. This network architecture is utilized for experiments on multi-to-single cross-domain generalization with fusion.
TABLE X: Architecture details for different modalities and fusion module.